The temperature of a 3.0L sample of gas is lowered from 350K to 275K. What will be the resulting volume of the gas?
Law being used: ___________ FORMULA WORK ANSWER
Given 1:__________
Given 2:__________
Given 3:__________
1
Expert's answer
2020-04-18T07:00:34-0400

Solution:

Pressure is constant.


Law being used: Charles' law


Given 1: __ V1 = 3.0 L __

Given 2: __ T1 = 350 K __

Given 3: __T2 = 275 K __


FORMULA:

V1 / T1 = V2 / T2


WORK:

V1*T2 = V2*T1

(3.0 L) * (275 K) = (V2) * (350 K)

V2 = (3.0 * 275) / 350 = 2.357 L

V2 = 2.357 L = 2.4 L


ANSWER: V2 = 2.4 L